OpenStudy (anonymous):

[6 - 6n] = 12. What is n?

OpenStudy (anonymous):

is this absolute value?

OpenStudy (anonymous):

if so, solve \[6-6n=12\] and \[6-6n=-12\] to get your two solutions

OpenStudy (anonymous):

{6-6n} = 12 6-6n =12 -6 -6 -6n = 6 /-6 /-6 n = -1